OpenStudy (anonymous):

Solve the system of linear equations below. y = -2x + 19 y = x + 7 answer choices below please help A)x = 4, y = 11 B)x = -12, y = -5 C)x = -4, y = 3 D)x = 8, y = 3

OpenStudy (anonymous):

combine alike things

OpenStudy (anonymous):

can you just tell me the answer?

OpenStudy (anonymous):

ok for a medal

OpenStudy (anonymous):

okay i gave u a medal i chouse best response

OpenStudy (anonymous):

the answer is a

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

did you get it correct

OpenStudy (anonymous):

idk im on a pretest for credit recovery

OpenStudy (anonymous):

why did you take my medal

OpenStudy (anonymous):

cause it was wrong

OpenStudy (anonymous):

ok my apologies

OpenStudy (anonymous):

its okay thanks for trying

OpenStudy (anonymous):

need any other help

OpenStudy (anonymous):

Janice wants to create a test containing 20 questions worth 50 points. If Janice creates questions worth either two points or four points, she can include ____ two-point questions and ______ four-point questions.

OpenStudy (anonymous):

what are the choices

OpenStudy (anonymous):

there are none i have to fill in the blanks ______ those are the blanks

OpenStudy (anonymous):

ok i think it is 15 for 2 and 5 for 4point questions

OpenStudy (anonymous):

An equation in the system is y = x + 1. What is the other equation in the system if the solution is only at (0, 1)? x − y = 1 2x + 2y = 4 x + y = 1 x + 2y = 2

OpenStudy (anonymous):

for the first answer choice its x − y = 1

OpenStudy (anonymous):

x - y